WAC Notes January 30 2020
|HOW TOs for shift crew|
|All Expert Contacts|
January 30 2020
Daily Meeting BlueJeans
No data today, the beam comes back at about 17:00, ATLis Procedures happened in the next hour, after that, we spent some time (~ half hour) to find the target tungsten collimator. Compton tuning (?) comes next, which didn't finish until 23:30 and was followed by BeamMod Commissioning. We start taking data at about 2:30 the next day (1.31).
Hall C requests IHWP switch twice, but we don't take any data, so Cameron changes the slug number back to 121 manually.
Non-Production Runs and Bad Runs
- FFB and orbit locks seem to be breaking, so there is no beam delivery for now